# Redistribution Through

### v2

@@ 5 Concluding remarks
Price discrimination has long been used as an instrument for redistribution. Many goods and services are provided at reduced prices for disadvantaged groups. For instance, students from low-income backgrounds often pay reduced tuition fees at universities, and it is common for the unemployed or the elderly to receive discounted prices at museums. The prevalence of such redistributive price discrimination indicates that personalized pricing is deemed socially acceptable as long as it is perceived to benefit those in need. This view is supported by experimental data on people’s attitudes to price discrimination (Wu et al., 2012) and is expressed in policy reports (White House Council of Economic Advisers, 2015; Bourreau and De Streel, 2018). Our redistributive segmentations satisfy this social demand by ensuring that price discrimination satisfies a certain degree of price progressivity. Hence, our results inform the current policy debate over the regulation of online personalized pricing. They give qualitative guidelines regarding market segmentations that prioritize consumers with low willingness to pay.
@@
Redistributive price discrimination is usually implemented by public and non-profit organizations (Le Grand, 1975). A concern is that imposing redistributive price discrimination on the private sector could be inefficient. Indeed, conditioning prices on group characteristics could harm profits and deter firms from entering the market (Woodcock, 2022). This paper studies how a private market with a profit-maximizing monopoly can be optimally segmented to achieve a redistributive goal. Redistributive segmentations have the twofold benefit of achieving economic efficiency (Lemma 1) and never harming the seller—Theorem 3 shows that redistributive segmentations can even improve the seller’s profit.
@@
One might suggest that conditioning market segmentation on additional buyer information (e.g., income or group characteristics), rather than only the willingness to pay, would better target redistribution. Nevertheless, as highlighted in Section 4, consumer data is usually collected and processed by private actors for which consumers’ willingness to pay is the payoff-relevant variable. It might also be impossible or illegal for regulators like competition authorities to directly access income data or protected characteristics. Hence, our optimal segmentations constitute a normative benchmark against which to evaluate welfare outcomes that could be achieved by regulating data intermediaries.
Another concern among policymakers is consumers’ privacy (Acquisti et al., 2016; Goldfarb and Que, 2023). The current approach mainly focuses on studying the incentives for consumers to reveal their preferences voluntarily (see Ichihashi, 2020; Hidir and Vellodi, 2020; Ali et al., 2023). Our redistributive segmentations are monotone Theorem 2.A, and thus separate rich and poor consumers in different market segments, sometimes strictly (Theorem 2.B). This separation might conflict with consumer privacy concerns, as it allows the seller to discern buyers’ types better. Studying market segmentations that prioritize poorer consumers’ welfare while satisfying privacy constraints (Strack and Yang, 2024) is a fruitful avenue for future research.
Appendix A Proof of Theorem 1
@@
We know from Lemma 1 that we can, without loss of generality, restrict attention to segmentations supported on $\Omega$. Hence, we treat a segmentation $\sigma$ as an element of $\mathbb{R}^{\Omega}$.
We first define the following set of $\mathcal{E}$ is the set of real functions $w$ defined on $\Omega$ that is supermodular and such that, for any $\theta\in\Theta$, the function $p\mapsto w(\theta,p)$ is non-increasing on $\{\theta_{1},\dots,\theta\}$.
=$\mathcal{E}=\Bigl{\{}v\in\mathbb{R}^{\Omega}\ \big{|}\ \exists w\in\mathcal{R}\text{ and }b\in\mathbb{R}\text{ such that},\forall(\theta,p)\in\Omega,\,v(\theta,p)=w(\theta,p)+b\Bigr{\}}.$
For any given aggregate market $\mu$, we define the order $\succsim_{\mathcal{E}}$ over the set $\Sigma_{\Omega}(\mu)$ in the analogous way we defined the order $\succsim_{\mathcal{R}}$. The following lemma allows us to work with the family of functions $\mathcal{E}$.
@@
We now prove the statement of the theorem for the order $\succsim_{\mathcal{E}}$. The proof follows similar steps as the proof of Theorem 1 in Meyer and Strulovici (2015). We first show that the set $\mathcal{E}$ is the dual cone of the set $\mathcal{T}$, as shown by the next lemma.
@@
$\Rightarrow$ This follows from the definition of $\mathcal{R}$ and therefore of $\mathcal{E}$.
$\Leftarrow$ Let $w$ be such that $\langle w,t\rangle\geq 0$ for every $t\in\mathcal{T}$. By contradiction, suppose there exists $(\theta,p),(\theta,p^{\prime})\in\Omega$ such that $p<p^{\prime}$ and $w(\theta,p)<w(\theta,p^{\prime})$. Take the downward buyer transfer $d\in\mathbb{R}^{\Omega}$ such that $d(\theta,p)=-d(\theta,p^{\prime})=1$, then $\langle w,d\rangle<0$, a contradiction. Similarly, suppose there exists $(\theta,p),(\theta^{\prime},p^{\prime})\in\Omega$ such that $\theta<\theta^{\prime}$, $p<p^{\prime}$, and $w(\theta,p)-w(\theta,p^{\prime})<w(\theta^{\prime},p)-w(\theta^{\prime},p^{\prime})$. In that case, take the redistributive buyer transfer $r\in\mathbb{R}^{\Omega}$ such that
$r(\theta,p)=r(\theta^{\prime},p^{\prime})=-r(\theta^{\prime},p)=-r(\theta,p^{\prime})=1,$
@@
Recall that the conic hull of $\mathcal{T}$ is the smallest convex cone containing $\mathcal{T}$ and equals the set of all positive linear combinations of finitely many elements of $\mathcal{T}$ (see Ok, 2007, p. 430):
@@
Note that for any $t\in\mathcal{T}$ and $\lambda_{t}\geq 0$, $\lambda_{t}t\in\mathcal{T}$, therefore we can define $\mathrm{cone}(\mathcal{T})$ as
@@
As $\mathrm{cone}(\mathcal{T})$ is closed, we can show in the next lemma that it is the dual cone of $\mathcal{E}$ (see Luenberger, 1969, p.215).
@@
$\Rightarrow$ Assume $\tau\in\mathrm{cone}(\mathcal{T})$. Hence, there exists $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that $\tau=\sum_{i=1}^{n}t_{i}$. Following Lemma 4 we have that, for any $w\in\mathcal{E}$,
@@
$\Leftarrow$ Assume $\upsilon\notin\mathrm{cone}(\mathcal{T})$. As $\mathrm{cone}(\mathcal{T})$ is closed, the separating hyperplane theorem implies that there exists $w\in\mathbb{R}^{\Omega}$ such that $\langle\upsilon,w\rangle<\langle\tau,w\rangle$ for every $\tau\in\mathrm{cone}(\mathcal{T})$. Since $\mathrm{cone}(\mathcal{T})$ is a cone, it must be that $\langle\tau,w\rangle\geq 0$ for any $\tau\in\mathrm{cone}(\mathcal{T})$ as otherwise there would exist some $\alpha>0$ such that $\langle\alpha\tau,w\rangle\leq\langle\upsilon,w\rangle$. Thus, Lemma 4 implies that $w\in\mathcal{E}$. Finally, since $0=\min_{\tau\in\mathrm{cone}(\mathcal{T})}\langle\tau,w\rangle$, we must have $\langle\upsilon,w\rangle<0$. ∎
We can now restate the theorem as follows: for any $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$, $\sigma\succsim_{\mathcal{E}}\sigma^{\prime}$ if and only if there exists $n\in\mathbb{N}^{*}$ and a sequence $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that:
@@
Equation 8 is equivalent to $\sigma-\sigma^{\prime}\in\mathrm{cone}(\mathcal{T})$, and by Lemma 5, this is equivalent to $\langle\sigma-\sigma^{\prime},w\rangle\geq 0$ for every $w\in\mathcal{E}$. ∎
Appendix B Proof of Proposition 1
Claim (i) was proven in the main text. We now prove claim (ii).
@@ Appendix B Proof of Proposition 1
For any set $A$ we let $-A=\bigl{\{}-a\ |\ a\in A\bigr{\}}$. For any two sets $A$ and $B$, we let $A+B\coloneqq\bigl{\{}a+b\ |\ a\in A,b\in B\bigr{\}}$ be the Minkowski sum of $A$ and $B$, and we let $A-B=A+(-B)$.
The set $\Sigma_{\Omega}(\mu)$ is a convex polytope. Let $\sigma\in\Sigma_{\Omega}(\mu)$ be $\succsim_{\mathcal{R}}$-maximal. Theorem 1 implies that for any $t\in\mathrm{cone}(\mathcal{T})$ such that $t\neq 0$, we must have $\sigma+t\notin\Sigma_{\Omega}(\mu)$. Therefore, $\bigl{[}\Sigma_{\Omega}(\mu)\cap\bigl{(}\{\sigma\}+\mathrm{cone}(\mathcal{T})\bigr{)}\bigr{]}=\{\sigma\}$. This implies that $\sigma$ is on the relative border of both $\Sigma_{\Omega}(\mu)$ and $\{\sigma\}+\mathrm{cone}(\mathcal{T})$. Given that these are closed convex sets, the separating hyperplane theorem implies that there exists $v\in\mathbb{R}^{\Omega}$ such that:
$\forall\sigma^{\prime}\in\Sigma_{\Omega}(\mu),\forall t\in\mathrm{cone}(\mathcal{T}),\quad\langle\sigma^{\prime},v\rangle\leq\langle\sigma,v\rangle\leq\langle\sigma+t,v\rangle.$
Thus, for any $t\in\mathrm{cone}(\mathcal{T})$,
$\langle\sigma,v\rangle\leq\langle\sigma+t,v\rangle\iff 0\leq\langle t,v\rangle,$
which, by Lemma 4, implies that $v\in\mathcal{E}$. Therefore, there exists $w\in\mathcal{R}$ and $b\in\mathbb{R}$ such that $v=w+b$. Hence, for any $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$,
$\langle\sigma^{\prime},v\rangle\leq\langle\sigma,v\rangle\iff\langle\sigma^{\prime},w+b\rangle\leq\langle\sigma,w+b\rangle\iff\langle\sigma^{\prime},w\rangle\leq\langle\sigma,w\rangle.$
Therefore, $\sigma$ is $\mathrm{P}^{\star}_{w,\mu}$-optimal. ∎
Appendix C Proof of Proposition 2
From Theorem 1, for any $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$, there exists no $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ such that $\sigma^{\prime}\succ_{\mathcal{R}}\sigma$ if and only if for any $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tau\neq 0$, $\sigma+\tau\notin\Sigma_{\Omega}(\mu)$. We first show the following equivalence.
@@ Appendix C Proof of Proposition 2
Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$. For any $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tau\neq 0$, $\sigma+\tau\notin\Sigma_{\Omega}(\mu)$ if and only if for any $t\in\mathcal{T}$ such that $t\neq 0$, $\sigma+t\notin\Sigma_{\Omega}(\mu)$.
@@
Necessity is straightforward. Let now $\sigma\in\Sigma_{\Omega}(\mu)$ be such that for any $t\in\mathcal{T}$ such that $t\neq 0$, $\sigma+t\notin\Sigma_{\Omega}(\mu)$. First, this means that $\sigma$ is on the relative border of $\Sigma_{\Omega}(\mu)$, as otherwise, for any neighbourhood around $\sigma$ included in $\Sigma_{\Omega}(\mu)$, there would exist sufficiently small transfers $t\in\mathcal{T}$ such that $\sigma+t$ is in this neighbourhood. Second, this also implies that $\sigma\neq\sigma^{F}$, where $\sigma^{F}$ is the first-degree price discrimination segmentation. That is, the segmentation that allocates each buyer to the segment where the price equals their type. This is because downward transfers could be performed on $\sigma^{F}$ without violating the obedience constraints.
@@
We now proceed by contradiction, by assuming that there exists $\tau\in\mathrm{cone}(\mathcal{T})$, $\tau\neq 0$, such that $\sigma^{\prime}\coloneqq\sigma+\tau\in\Sigma_{\Omega}(\mu)$, that is, $\sigma^{\prime}\succ_{\mathcal{R}}\sigma$. As in Appendix B, we let $-A=\bigl{\{}-a\ |\ a\in A\bigr{\}}$ for any set $A$, and $A+B\coloneqq\bigl{\{}a+b\ |\ a\in A,b\in B\bigr{\}}$ and $A-B=A+(-B)$ for any two sets $A$ and $B$. We first prove that $\Sigma_{\Omega}(\mu)\subset\bigl{(}\{\sigma\}+\mathrm{cone}(\mathcal{T})\bigr{)}$, which is equivalent to $\bigl{(}\Sigma_{\Omega}(\mu)-\{\sigma\}\bigr{)}\subset\mathrm{cone}(\mathcal{T})$. Given that $\sigma^{\prime}=\sigma+\tau\in\Sigma_{\Omega}(\mu)$, for some $\tau\neq 0$ and $\tau\notin\mathcal{T}$ this means that $\bigl{(}\Sigma_{\Omega}(\mu)-\{\sigma\}\bigr{)}\,\cap\,\operatorname{int}\bigl{(}\mathrm{cone}(\mathcal{T})\bigr{)}\neq\varnothing$. Now suppose by contradiction that $\bigl{(}\Sigma_{\Omega}(\mu)-\{\sigma\}\bigr{)}\not\subset\mathrm{cone}(\mathcal{T})$. As $\Sigma_{\Omega}(\mu)$ and $\mathrm{cone}(\mathcal{T})$ are closed and convex sets, there exists $x\in\Sigma_{\Omega}(\mu)$ that is on relative border of $\mathrm{cone}(\mathcal{T})$. This means that there exists $\tilde{\sigma}\in\Sigma_{\Omega}(\mu)$ such that $\tilde{\sigma}=\sigma+t$ for some $t\in\mathcal{T}$ (as the borders of $\mathrm{cone}(\mathcal{T})$ are elements of $\mathcal{T}$), which contradicts our premise. Hence we proved that $\Sigma_{\Omega}(\mu)-\{\sigma\}\subset\mathrm{cone}(\mathcal{T})$; which implies that for any $\tilde{\sigma}\in\Sigma_{\Omega}(\mu)$, there exists $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tilde{\sigma}=\sigma+\tau$. To see that this is a contradiction, consider $\sigma^{F}$. We know that $\sigma^{F}\in\Sigma_{\Omega}(\mu)$ and $\sigma^{F}\neq\sigma$, but $\sigma^{F}$ cannot be reached from $\sigma$ by a finite sequence of transfers in $\mathcal{T}$; that is, there exists no $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\sigma^{F}=\sigma+\tau$. ∎
@@
Equipped with Lemma 6, we can now proceed to the proof of Proposition 2 by proving it is equivalent that a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ satisfies claims (i) and (ii) in the proposition, and that there exists no $t\in\mathcal{T}$ such that $t\neq 0$ and $\sigma+t\in\Sigma_{\Omega}(\mu)$. For any $\mu\in\Delta(\Theta)$, $\sigma\in\Sigma_{\Omega}(\mu)$ and $p\in\operatorname{supp}(\sigma_{P})$, define:
@@
Proof of Proposition 2.
$\Rightarrow$ Suppose that that there exists no $t\in\mathcal{T}$ such that $t\neq 0$ and $\sigma+t\in\Sigma_{\Omega}(\mu)$. Note first that downward and redistributive buyer transfers do not change the marginal distribution over types. Hence, for any $\sigma^{\prime}$ obtained from $\sigma$ by such transfers, we can restrict our attention to the case where $\sigma^{\prime}\notin\Sigma_{\Omega}(\mu)$ is caused by $\sigma^{\prime}$ violating Ob. We first prove that claim (i) in Proposition 2 is satisfied. Denote $\bar{p}=\max\operatorname{supp}(\sigma_{P}))$ and $\bar{\theta}=\max\bigl{\{}\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,\bar{p})\bigr{)}\bigr{\}}$, and let $p\in\operatorname{supp}(\sigma_{P})$ be such that $p<\bar{p}$. As there exists no $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ that is obtained from $\sigma$ by a downward buyer transfer, this means that there exists $p<q\leq\bar{\theta}$ such that $q\in\mathcal{I}_{\sigma}(p)$, as otherwise, we could construct $\sigma^{\prime}$ from $\sigma$ by a downward buyer transfer of mass $\varepsilon>0$ from $(\bar{\theta},\bar{p})$ to $(\bar{\theta},p)$ such that $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ for $\varepsilon$ sufficiently small. This implies that $|\mathcal{I}_{\sigma}(p)|\geq 2$.
We now prove claim (ii). Let $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$. By contradiction, assume there exists $q\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}$ such that $q\geq p^{\prime}$ and that $q\notin\mathcal{I}_{\sigma}(p^{\prime})$. This implies that $q>p^{\prime}$, so that the greatest type in $\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}$ that is strictly lower than $q$, defined by

### v5

@@ 5. Concluding Remarks
Price discrimination has long been used as an instrument for redistribution. Many goods and services are provided at reduced prices for disadvantaged groups. For instance, students from low-income backgrounds often pay reduced tuition fees at universities, and it is common for the unemployed or the elderly to receive discounted prices at museums. The prevalence of such redistributive price discrimination indicates that personalized pricing is deemed socially acceptable as long as it is perceived to benefit those in need. This view is supported by experimental data on people’s attitudes to price discrimination (wu) and is expressed in policy reports (WhiteHouse2015; bourreau2018regulation). Our redistributive segmentations satisfy this social demand by ensuring that price discrimination satisfies a certain degree of price progressivity. Hence, our results inform the current policy debate over the regulation of online personalized pricing. They give qualitative guidelines regarding market segmentations that prioritize consumers with low willingness to pay.
@@
Redistributive price discrimination is usually implemented by public and non-profit organizations (legrand; steinberg). A concern is that imposing redistributive price discrimination on the private sector could be inefficient. This paper studies how a private market with a profit-maximizing monopoly can be optimally segmented to achieve a redistributive goal. Redistributive segmentations have the twofold benefit of achieving economic efficiency (Lemma˜1) and never harming the seller—Proposition˜3 shows that redistributive segmentations can even improve the seller’s profit.
@@
One might suggest that conditioning market segmentation on additional buyer information (e.g., income or group characteristics), rather than only the willingness to pay, would better target redistribution. Nevertheless, as highlighted in Section˜4, consumer data is usually collected and processed by private actors for which consumers’ willingness to pay is the payoff-relevant variable. It might also be impossible or illegal for regulators like competition authorities to directly access income data or protected characteristics. Hence, our optimal segmentations constitute a normative benchmark against which to evaluate welfare outcomes that could be achieved by regulating data intermediaries.
Appendices
Appendix A Proof of Proposition˜1
@@
We know from Lemma˜1 that we can, without loss of generality, restrict attention to segmentations supported on $\Omega$. Hence, we treat a segmentation $\sigma$ as an element of $\mathbb{R}^{\Omega}$.
[…]
$\mathcal{E}=\Bigl\{v\in\mathbb{R}^{\Omega}\ \big|\ \exists w\in\mathcal{R}\text{ and }b\in\mathbb{R}\text{ such that},\forall(\theta,p)\in\Omega,\,v(\theta,p)=w(\theta,p)+b\Bigr\}.$
[…]
@@
[…]
@@
Necessity follows from the definition of $\mathcal{R}$ and therefore of $\mathcal{E}$. Let us prove sufficiency. Let $w$ be such that $\langle w,t\rangle\geq 0$ for every $t\in\mathcal{T}$. By way of contradiction, assume that there exists $(\theta,p),(\theta,p^{\prime})\in\Omega$ such that $p<p^{\prime}$ and $w(\theta,p)<w(\theta,p^{\prime})$. Consider the downward buyer transfer $d\in\mathbb{R}^{\Omega}$ such that $d(\theta,p)=-d(\theta,p^{\prime})=1$. Then $\langle w,d\rangle<0$, a contradiction. Assume now that there exists $(\theta,p),(\theta^{\prime},p^{\prime})\in\Omega$ such that $\theta<\theta^{\prime}$, $p<p^{\prime}$, and $w(\theta,p)-w(\theta,p^{\prime})<w(\theta^{\prime},p)-w(\theta^{\prime},p^{\prime})$. Consider the redistributive buyer transfer $r\in\mathbb{R}^{\Omega}$ such that
$r(\theta,p)=r(\theta^{\prime},p^{\prime})=-r(\theta^{\prime},p)=-r(\theta,p^{\prime})=1.$
@@
Let $\mathrm{cone}(\mathcal{T})$ denote the conic hull of $\mathcal{T}$, the smallest convex cone containing $\mathcal{T}$, equal to the set of all positive linear combinations of finitely many elements of $\mathcal{T}$ (see ok2007real, p. 430):
@@
Note that for any $t\in\mathcal{T}$ and $\lambda_{t}\geq 0$, $\lambda_{t}t\in\mathcal{T}$. Therefore,
@@
As $\mathrm{cone}(\mathcal{T})$ is closed, we can show in the next lemma that it is the dual cone of $\mathcal{E}$ (see luenberger1969optimization, p.215).
@@
Assume that $\tau\in\mathrm{cone}(\mathcal{T})$. Hence, there exists $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that $\tau=\sum_{i=1}^{n}t_{i}$. Following Lemma˜6 we have that, for any $w\in\mathcal{E}$,
@@
Assume now that $\upsilon\notin\mathrm{cone}(\mathcal{T})$. As $\mathrm{cone}(\mathcal{T})$ is closed, the Separating Hyperplane Theorem implies that there exists $w\in\mathbb{R}^{\Omega}$ such that $\langle\upsilon,w\rangle<\langle\tau,w\rangle$ for every $\tau\in\mathrm{cone}(\mathcal{T})$. Since $\mathrm{cone}(\mathcal{T})$ is a cone, it must be that $\langle\tau,w\rangle\geq 0$ for any $\tau\in\mathrm{cone}(\mathcal{T})$ as otherwise there would exist some $\alpha>0$ such that $\langle\alpha\tau,w\rangle\leq\langle\upsilon,w\rangle$. Thus, Lemma˜6 implies that $w\in\mathcal{E}$. Finally, since $0=\min_{\tau\in\mathrm{cone}(\mathcal{T})}\langle\tau,w\rangle$, we must have $\langle\upsilon,w\rangle<0$. ∎
We can now restate the theorem as follows: for any $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$, $\sigma\succsim_{\mathcal{E}}\sigma^{\prime}$ if and only if there exists $n\in\mathbb{N}^{*}$ and a sequence $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that:
@@
Equation˜5 is equivalent to $\sigma-\sigma^{\prime}\in\mathrm{cone}(\mathcal{T})$, and by Lemma˜7, this is equivalent to $\langle\sigma-\sigma^{\prime},w\rangle\geq 0$ for every $w\in\mathcal{E}$. ∎
Appendix B Proof of Lemma˜2
@@
[…]
We now prove claim (ii). For any set $A$ we let $A=\bigl\{-a\ |\ a\in A\bigr\}$. For any two sets $A$ and $B$, we let $A+B\coloneqq\bigl\{a+b\ |\ a\in A,b\in B\bigr\}$ be the Minkowski sum of $A$ and $B$, and we let AB=A+(B)A-B=A+(-B).
[…]
Given that these are closed convex sets, the separating hyperplane theorem implies that there exists $v\in\mathbb{R}^{\Omega}$ such that:
$\forall\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu),\forall t\in\mathrm{cone}(\mathcal{T}),\quad\langle\sigma^{\prime},v\rangle\leq\langle\sigma,v\rangle\leq\langle\sigma+t,v\rangle.$
Furthermore, for any $t\in\mathrm{cone}(\mathcal{T})$, $\langle\sigma,v\rangle\leq\langle\sigma+t,v\rangle\iff 0\leq\langle t,v\rangle$, which, by Lemma˜6, implies that $v\in\mathcal{E}$. Therefore, there exists $w\in\mathcal{R}$ and $b\in\mathbb{R}$ such that $v=w+b$. Hence, for any $\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$,
$\langle\sigma^{\prime},v\rangle\leq\langle\sigma,v\rangle$ $\iff\langle\sigma^{\prime},w+b\rangle\leq\langle\sigma,w+b\rangle$
$\iff\langle\sigma^{\prime},w\rangle\leq\langle\sigma,w\rangle,$
which proves that $\sigma$ is $\mathrm{P}^{\star}_{w,\mu}$-optimal. ∎
Appendix C Proof of Lemma˜3
From Proposition˜1, for any $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$, there exists no $\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$ such that $\sigma^{\prime}\succ_{\mathcal{R}}\sigma$ if and only if for any $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tau\neq 0$, $\sigma+\tau\notin\Sigma^{\mathrm{eff}}(\mu)$. We first show the following equivalence.
@@ Appendix C Proof of Lemma˜3
Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$. For any $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tau\neq 0$, $\sigma+\tau\notin\Sigma^{\mathrm{eff}}(\mu)$ if and only if for any $t\in\mathcal{T}$ such that $t\neq 0$, $\sigma+t\notin\Sigma^{\mathrm{eff}}(\mu)$.
@@
Necessity is straightforward. Let now $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ be such that for any $t\in\mathcal{T}$ such that $t\neq 0$, $\sigma+t\notin\Sigma^{\mathrm{eff}}(\mu)$. First, this means that $\sigma$ is on the relative border of $\Sigma^{\mathrm{eff}}(\mu)$, as otherwise, for any neighborhood around $\sigma$ included in $\Sigma^{\mathrm{eff}}(\mu)$, there would exist sufficiently small transfers $t\in\mathcal{T}$ such that $\sigma+t$ is in this neighborhood. Second, this also implies that $\sigma\neq\sigma^{F}$, where $\sigma^{F}$ is the first-degree price discrimination segmentation. That is, the segmentation supported on the line $p=\theta$. This is because downward transfers could be performed on $\sigma^{F}$ without violating the obedience constraints.
@@
We now proceed by contradiction, by assuming that there exists $\tau\in\mathrm{cone}(\mathcal{T})$, $\tau\neq 0$, such that $\sigma^{\prime}\coloneqq\sigma+\tau\in\Sigma^{\mathrm{eff}}(\mu)$, that is, $\sigma^{\prime}\succ_{\mathcal{R}}\sigma$. As in Appendix˜B, we let $A=\bigl\{-a\ |\ a\in A\bigr\}$ for any set $A$, and $A+B\coloneqq\bigl\{a+b\ |\ a\in A,b\in B\bigr\}$ and AB=A+(B)A-B=A+(-B) for any two sets $A$ and $B$. We first prove that $\Sigma^{\mathrm{eff}}(\mu)\subset\bigl(\{\sigma\}+\mathrm{cone}(\mathcal{T})\bigr)$, which is equivalent to $\bigl(\Sigma^{\mathrm{eff}}(\mu)-\{\sigma\}\bigr)\subset\mathrm{cone}(\mathcal{T})$. Given that $\sigma^{\prime}=\sigma+\tau\in\Sigma^{\mathrm{eff}}(\mu)$, for some $\tau\neq 0$ and $\tau\notin\mathcal{T}$ this means that $\bigl(\Sigma^{\mathrm{eff}}(\mu)-\{\sigma\}\bigr)\,\cap\,\operatorname{int}\bigl(\mathrm{cone}(\mathcal{T})\bigr)\neq\varnothing$. Now suppose by contradiction that $\bigl(\Sigma^{\mathrm{eff}}(\mu)-\{\sigma\}\bigr)\not\subset\mathrm{cone}(\mathcal{T})$. As $\Sigma^{\mathrm{eff}}(\mu)$ and $\mathrm{cone}(\mathcal{T})$ are closed and convex sets, there exists $x\in\Sigma^{\mathrm{eff}}(\mu)$ that is on relative border of $\mathrm{cone}(\mathcal{T})$. This means that there exists $\tilde{\sigma}\in\Sigma^{\mathrm{eff}}(\mu)$ such that $\tilde{\sigma}=\sigma+t$ for some $t\in\mathcal{T}$ (as the borders of $\mathrm{cone}(\mathcal{T})$ are elements of $\mathcal{T}$), which contradicts our premise. Hence we proved that $\Sigma^{\mathrm{eff}}(\mu)-\{\sigma\}\subset\mathrm{cone}(\mathcal{T})$; which implies that for any $\tilde{\sigma}\in\Sigma^{\mathrm{eff}}(\mu)$, there exists $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tilde{\sigma}=\sigma+\tau$. To see that this is a contradiction, consider $\sigma^{F}$. We know that $\sigma^{F}\in\Sigma^{\mathrm{eff}}(\mu)$ and $\sigma^{F}\neq\sigma$, but $\sigma^{F}$ cannot be reached from $\sigma$ by a finite sequence of transfers in $\mathcal{T}$; that is, there exists no $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\sigma^{F}=\sigma+\tau$. ∎
@@
Equipped with Lemma˜8, we can now proceed to the proof of Lemma˜3 by proving it is equivalent that a segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ is saturated and that there exists no $t\in\mathcal{T}$ such that $t\neq 0$ and $\sigma+t\in\Sigma^{\mathrm{eff}}(\mu)$. For any $\mu\in\Delta(\Theta)$, $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ and $p\in\operatorname{supp}(\sigma_{P})$, define:
@@
Proof of Lemma˜3.
[…]
We now prove part (b) in Definition˜1. Let $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$. By contradiction, assume there exists $q\in\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,p)\bigr)$ such that $q\geq p^{\prime}$ and that $q\notin\mathcal{I}_{\sigma}(p^{\prime})$. This implies that $q>p^{\prime}$, so that the greatest type in $\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,p)\bigr)$ that is strictly lower than $q$, defined by